What is the mass of 1.56 x 1023 atoms of aluminum?
1
Expert's answer
2020-09-15T13:49:59-0400

The molecular weight of Aluminum "=26.98g\/mol"

The mass of Aluminum is unknown hence "(x)"

Atoms of Aluminum"=1.56\times 10^{23}"

In determination of the unknown value, we convert from Atoms"\to" Moles"\to" Mass

Atoms"\to" Moles

Moles of Aluminum( using Avogadros constant) is also unknown"(y)"

"6.022\times10^{23}Al\over1molAl" "\times ymol Al=1.56\times10^{23}AtomsAl"

"y mol Al=""1.56\times 10^{23} \over 6.022\times 10^{23}" "=0.259mol Al"

Moles"\to" Mass

"xgAl\times" "1molAl\over 26.98gAl" "=0.259 mol Al" .

"xgAl=(0.259 mol Al\times 26.98g/mol) =6.98g Al"

The Mass is thus "6.98g"